OpenStudy (anonymous):

Can you solve this analogy 6:9 :: 64:?

TheSmartOne (thesmartone):

Is your question 6:9 :: 6:4 ?`

OpenStudy (anonymous):

No it's "64"

TheSmartOne (thesmartone):

Oh I get it...

TheSmartOne (thesmartone):

a:b is noting but a/b

OpenStudy (anonymous):

Its a "Complete the analogy" kind of a question! Ya... I also did that and got 96 . But the correct answer according to the book (without explanation, of course!) is 25.

TheSmartOne (thesmartone):

So \[\frac{ 6 }{ 9 }=\frac{ 64 }{ ? }\]

TheSmartOne (thesmartone):

Well... This is the first time I hvae seen this question in math...

TheSmartOne (thesmartone):

@mathslover @jhonyy9 @Loser66 @paki Do you get this question?

OpenStudy (anonymous):

Maybe I could repost it in a Logic subject section, if there is one on this website.

TheSmartOne (thesmartone):

There is no logic website... You can keep it open in Math...

mathslover (mathslover):

Well, its a question of proportion.

mathslover (mathslover):

6:9 :: 64:? \(\cfrac{6}{9} = \cfrac{64}{?}\) \(\cfrac{6}{9} \times 9 = \cfrac{64}{?} \times 9 \\ \cfrac{6}{\cancel{9}} \times \cancel{9} = \cfrac{(64 \times 9 )}{?} \) So, \( 6 = \cfrac{(64 \times 9)}{?}\)

mathslover (mathslover):

Further, multiply both sides by "?" \(6 \times ? = 64 \times 9 \) Divide both sides by 6 \(? = \cfrac{64 \times 9 }{ 6}\)

jhonyy9 (jhonyy9):

why you dont simplifie 6/9 = 2/3 ???

mathslover (mathslover):

You can, but I preferred to go through the conventional method @jhonyy9 !

jhonyy9 (jhonyy9):

ok ! but i think the result will be the same

mathslover (mathslover):

Yep! Result will be same in both cases.

OpenStudy (anonymous):

Yup, I did the same. The answer , I think should be 96. Maybe the answer in the book is wrong. Thanks anyway! @mathslover @jhonyy9 @TheSmartOne

mathslover (mathslover):

You're welcome Mttal :)

TheSmartOne (thesmartone):

No problem...